# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@ 4 LLM Debugging with Sample-based Explanation
In this section, we discuss sample-based explanation strategies for LLMs, which aim to trace back the answers generated by LLMs to specific training samples (i.e., documents) or document fragments in the corpora. The utility of sample-based explanations for LLMs is multifaceted. First, tracing back the predictions of LLM to the training samples can provide evidence for the generation results, which facilitates model debugging in cases of errors and increases the trustworthiness of the model from users when the outcomes are accurate. In addition, it can also help researchers understand how LLMs generalize from training samples. If the outputs of LLMs can be traced back to exact subsequences directly spliced from the training data, it might suggest that the LLM is simply memorizing the data. In contrast, if the generation results and the influencing training samples are abstractly related, it could indicate that LLMs can understand the concepts and generate responses by reasoning from input prompts.
In this section, we start by systematically reviewing traditional sample-based explanation strategies, including gradient-based methods and embedding-based methods, as well as some preliminary explorations to generalize them to LLMs. We then analyze the challenges associated with generalizing the above strategies to LLMs with unique transformer structures and unprecedented numbers of parameters. Finally, we discuss the insights to address the challenges, as well as open challenges worthy of further investigation.
4.1 Literature Review of Sample-based Explanation
In this section, we denote the input space and output space as $\mathcal{X}$ and $\mathcal{Y}$, respectively. In the context of large language models (LLMs), $\mathcal{X}$ is the space of token sequences known as the prompts, and $\mathcal{Y}$ could be the space of discrete labels in classification tasks or the space of token sequences as output in generation tasks***Here, we should note that, in the pretraining phase of LLM with language modeling, the model either predicts masked tokens in $x_{i}$ (i.e., masked language modeling), or autoregressively predicts the next token (i.e., causal language modeling) in $x_{i}$. Therefore, in some works, $y_{i}$ is omitted, and only $x_{i}$ is included for discussions.. Accordingly, we have a training dataset $\mathcal{D}_{train}=\{z_{i}=(x_{i},y_{i})\}_{i=1}^{N}$ with $N$ samples drawn from the joint space $\mathcal{X}\times\mathcal{Y}$, on which an LLM model $f_{\theta}$ is trained with pretrained parameters $\hat{\theta}\in\mathbb{R}^{P}$. We also have a test sample $z=(x,y)$ of interest, where we want to explain the generation of $y$ from $x$ based on training samples in $\mathcal{D}_{train}$ (which can be viewed as the information source). The goal of sample-based explanation is to measure the influence of a training sample $z_{i}\in\mathcal{D}_{train}$ or a certain segment within $z_{i}$, such that the generation of LLMs can be well-explained and backed up by the selected training samples.
4.1.1 Influence Function-based Methods
@@
One strategy to quantify the influence of a training sample $z_{i}$ in the dataset $\mathcal{D}_{train}$ to a test sample $z$ is through the influence function (Koh & Liang, 2017; Han et al., 2020). It measures the change of the prediction loss $\mathcal{L}\left(z,\theta\right)$ for the test sample $z$, when the training sample $z_{i}$ undergoes a hypothetical modification in the dataset $\mathcal{D}_{train}$ during model training. This modification results in an altered set of optimal model parameters, denoted as $\hat{\theta}_{-z_{i}}$. The most common modification of a training sample is to remove it from the dataset, where the influence of the removal of a training sample $z_{i}$ on the loss at test sample $z$ can be computed as follows:
$\mathcal{I}(z_{i},z)=-\nabla_{\theta}\mathcal{L}(z,\hat{\theta})^{\top}\mathbf{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\mathcal{L}(z_{i},\hat{\theta}),$ (1)
@@
where $\nabla_{\theta}\mathcal{L}(z,\hat{\theta})$ is the gradient of the loss function $\mathcal{L}$ on the test sample $z$ evaluated at the optimal parameters $\hat{\theta}$, and $\mathbf{H}_{\hat{\theta}}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}^{2}\mathcal{L}(z_{i},\hat{\theta})$ denotes the Hessian matrix of the LLM model at parameter $\hat{\theta}$. If we denote the number of parameters in $\hat{\theta}$ as $P$, the naïve inversion of the Hessian matrix $\mathbf{H}_{\hat{\theta}}$ leads to $\mathcal{O}(NP^{2}+P^{3})$ time complexity and $\mathcal{O}(P^{2})$ space complexity (Schioppa et al., 2022), which is clearly infeasible for large models. To improve efficiency, Koh & Liang (2017) adopt an iterative approximation process, i.e., LiSSA (Linear time Stochastic Second-Order Algorithm), to calculate the Hessian-Vector Product (HVP) in Eq. (1), where the memory complexity can be reduced to $\mathcal{O}(P)$ and time complexity to $\mathcal{O}(NPr)$ ($r$ is the number of iterations). To further reduce the complexity, Pruthi et al. (2020) propose an alternative to Eq. (1), i.e., TracIn, which measures the influence of $z_{i}$ on $z$ by calculating the total reduction of the loss on $z$ whenever $z_{i}$ is included in the minibatch during model training. The TracIn measurement can be formulated as follows:
@@
$\mathcal{I}_{\operatorname{TracIn}}\left(z_{i},z\right)=\sum_{t:z_{i}\in\mathcal{B}_{t}}\mathcal{L}\left(z,\theta_{t}\right)-\mathcal{L}\left(z,\theta_{t+1}\right)\approx\frac{1}{b}\sum_{t:z_{i}\in B_{t}}\eta_{t}\nabla_{\theta}\mathcal{L}\left(z_{i},\theta_{t}\right)\cdot\nabla_{\theta}\mathcal{L}\left(z,\theta_{t}\right),$ (2)
@@
where $\mathcal{B}_{t}$ is the $t$-th mini-batch fed into the model during training, $\theta_{t}$ is the parameter checkpoint at the $t$-th step, $\eta_{t}$ is the step size, and $b$ is the size of the mini-batch. According to the above equation, TracIn only leverages gradient terms, where Hessian $\mathbf{H}_{\hat{\theta}}$ is removed from the influence measurement. This substantially improves the efficiency. However, such complexity is still prohibitive for large models from both the computational and memory perspectives. In addition, TracIn can only estimate the influence of adding/removing the sample to the loss, where variants of the vanilla influence function defined in Eq. (1) can measure the influence of other modifications of the training sample $z_{i}$, such as perturbation (e.g., masking out a segment of a document $x_{i}$). To adapt the vanilla influence function of Eq. (1) to explain transformers, Schioppa et al. (2022) propose to use Alnordi iteration (Arnoldi, 1951) to find the dominant eigenvalues and eigenvectors of the Hessian matrix on randomly sampled subsets $\mathcal{D}_{sub}$, with $|\mathcal{D}_{sub}|\ll|\mathcal{D}_{train}|$. In such a case, the diagonalized Hessian can be cheaply cached and inverted, where the computational and memory complexity can be substantially reduced. Previous work mainly focuses on reducing the complexity of calculating the influence of a single training sample. Observing that finding the most influential training sample on $z$ needs to iterate Eq. (1) overall $N$ training samples, Guo et al. (2021) propose to use fast KNN to pre-filter a small subset of influence-worthy data points from $\mathcal{D}_{train}$ as candidates to explain small pretrained language models, whereas Han & Tsvetkov (2022) propose to iteratively find a small subset $\mathcal{D}_{sub}\subset\mathcal{D}_{train}$ whose gradient is the most similar to that of the downstream task examples. Recently, Grosse et al. (2023) propose to use the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) approximation to scale influence functions to LLMs up to 52B parameters. For adaptation, only influences mediated by the multi-layer perceptron (MLP) layers are considered, whereas weights from the self-attention layers are fixed, as previous work has demonstrated that knowledge is mainly encoded in the MLP layers (Wang et al., 2023d). In addition, based on the assumption that weights from different MLP layers are independent, the EK-FAC approximated influence can be formulated as the sum of influences mediated by each layer:
$\mathcal{I}_{\text{EKFAC}}(z_{i},z)=\sum_{l}\nabla_{\theta^{(l)}}\mathcal{L}(z,\hat{\theta})^{\top}(\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}+\lambda^{(l)}\mathbf{I})^{-1}\nabla_{\theta^{(l)}}\mathcal{L}(z_{i},\hat{\theta}),$ (3)
@@
where $\theta^{(l)}$ denotes the weights of the $l$-th MLP layer, and $\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}$ is the EK-FAC approximated Gauss-Newton Hessian for $\theta^{(l)}$. Since the inversion of $L$ small ${K_{l}\times K_{l}}$ matrices (i.e., $\mathcal{O}(L\times K_{l}^{3})$) is substantially more efficient than the inversion of a large $L{K_{l}\times LK_{l}}$ matrix (i.e., $\mathcal{O}{((LK_{l}})^{3})$), $\mathcal{I}_{\text{EKFAC}}$ can be adaptable to very large models, let alone the HVP can be further simplified by utilizing the EK-FAC properties.
Recently, influence function-based explanation has been used in efficient finetuning of LLMs, where influence function has been used to select a small subset of training samples given few-shot validation samples for a specific downstream task, where the training overhead can be substantially improved (Xia et al., 2024).
4.1.2 Embedding-based Methods
@@
Another strategy for sample-based explanation involves leveraging the hidden representations within the transformer architecture, which is recognized for encoding high-level semantics from textual data, to calculate the semantic similarity between $z$ and $z_{i}$. The similarity can also be used to measure the influence of $z_{i}$ on $z$ as explanations (Rajani et al., 2019). Specifically, Akyurek et al. (2022) propose to represent the training sample $z_{i}$ and test sample $z$ by concatenating the input and output as $z^{cat}_{i}=[x_{i}||y_{i}]$, $z^{cat}=[x||y]$. The concatenation is feasible for generation tasks where the output $y$ lies in the same token sequence space as the input prompt $x$. The similarity between $z_{i}$ and $z$ can then be calculated as follows:
$\mathcal{I}_{\text{emb}}\left(z_{i},z\right)=\frac{f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\cdot f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)}{\left\|f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\right\|\left\|f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)\right\|},$ (4)
@@
where $f^{(l)}_{\hat{\theta}}$ is the sub-network that outputs the $l$-th layer intermediate activation of the pretrained LLM $f_{\hat{\theta}}$. The Eq. (4) has a similar form as the vanilla influence function defined in Eq. (1) as well as its TracIn alternative defined in Eq. (2), which assigns a score $\mathcal{I}$ for the explainee $z$ for each training sample $z_{i}$ in the dataset $\mathcal{D}_{train}$ as the explanation confidence of the sample $z_{i}$.
Compared with the influence function methods introduced in the previous part, embedding-based methods are computationally efficient, as for each explainee $z$, the explanation score from a training sample $z_{i}$ requires only one forward pass of the transformer network. In addition, the calculation can be easily paralleled for different training samples. However, the disadvantage is also evident: These methods lack a theoretical foundation and may fail to identify important training samples that may not be semantically similar to the test sample. Consider the following toy example: Training samples $z_{i}$ = (“1+1=”, “2”) and $z_{j}$ = (“2+2=”, “4”) make the LLM gain the ability to conduct arithmetic calculation, which explains why prompting the model with $x$ = “100+100” gives the results $y$=“200”. However, the embeddings between the test sample $z$ and the two training samples $z_{i}$ and $z_{j}$ can be very different when calculated via Eq. (4(Akyurek et al., 2022). Therefore, embedding-based methods may not be able to faithfully find the training samples where the explanations require generalization ability beyond semantic similarity.
4.2 Case Study: EK-FAC-based Influence Estimation
@@
In this part, we implement the EK-FAC-approximated influence function proposed in Grosse et al. (2023), and verify its scalability and effectiveness on LLMs with billions of parameters, including GPT2-1.5B (Radford et al., 2019), LLaMA2-7B (Touvron et al., 2023b), Mistral-7B (Jiang et al., 2023), and LLaMA2-13B.
4.2.1 Experimental Design
@@
We use the SciFact dataset (Wadden et al., 2020) as the corpora, which contains the abstract of 5,183 papers from the domain of basic science and medicine. The explainee LLMs are obtained by finetuning the pretrained LLMs for 20,000 iterations, where AdamW (Loshchilov & Hutter, 2018) is used as the optimizer, and the learning rate and weight decay are set to 1e-5, 1e-3, respectively. Then, we use 500 samples from the corpora to estimate the (i) uncentered covariance matrices of the activations and pre-activation pseudo-gradients $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$, and (ii) the variances of the projected pseudo-gradient $\bm{\Lambda}^{(l)}$ for each selected dense layer $l$, and cache them on the hard disk (details see Eqs. (16) and (20) in Grosse et al. (2023)). We select the c_fc layer for GPT2-1.5B, and gate_proj layer for LLaMA2-7B, Mistral-7B, and LLaMA2-13B***All the implementation and layer names are based on the huggingface transformers, where the details can be found in https://huggingface.co/docs/transformers/en/index..
For evaluation, we randomly select 200 samples from the corpora to construct the test set, which we name SciFact-Inf. Specifically, for the $j$-th selected sample $z_{j}=(x_{j},x_{j})$ (here $y_{j}=x_{j}$ as label equals the input in language modeling), we use the first three sentences in $x_{j}$, i.e., $\hat{x}_{j}$, to generate a completion $\hat{y}_{j}$ with the finetuned LLM (here, $\hat{y}_{j}$ does not equal the remaining sentences in $x_{j}$), and we aim to explain the generation of $\hat{y}_{i}$ from $\hat{x}_{j}$ with the finetuned LLM with the training samples via EK-FAC approximated influence scores defined in Eq. (3). Ideally, the $j$-th training sample $z_{j}$ itself should be the most influential sample w.r.t. the generation of $\hat{y}_{j}$ for test sample $\hat{z}_{j}$, which facilitates quantitative analysis of the effectiveness of Eq. (3).
In our implementation, for each test sample $\hat{z}_{j}$, we first calculate the EK-FAC approximated HVP part of the influence $\mathcal{I}_{\text{EKFAC}}(z_{i},\hat{z}_{j})$, i.e., $\sum_{l}\nabla_{\theta^{(l)}}\mathcal{L}(\hat{z}_{j},\hat{\theta})^{\top}(\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}+\lambda^{(l)}\mathbf{I})^{-1}$, which is shared for all training samples $z_{i}$. Specifically, we record the layer-wise gradient $\nabla_{\theta^{(l)}}\mathcal{L}(\hat{z}_{j},\hat{\theta})$ and calculate the HVP with the cached $\mathbf{Q}_{A}^{(l)}$, $\mathbf{Q}_{S}^{(l)}$ as Eq. (21) in Grosse et al. (2023). We then go through candidate training samples (1 positive and 99 negative), calculate the gradient $\nabla_{\theta^{(l)}}\mathcal{L}(z_{i},\hat{\theta})$, and take inner-product with the approximate HVP as the layer-wise influence. Finally, the layer-wise influences are summed up as Eq. (3) as the total influence $\mathcal{I}_{\text{EKFAC}}(z_{i},\hat{z}_{j})$. We rank the influence and calculate the top-$K$ hit rate of the positive training sample.
4.2.2 Results and Analysis

### v2

@@ 4. LLM Debugging with Sample-based Explanation
In this section, we discuss sample-based explanation strategies for LLMs, which aim to trace back the answers generated by LLMs to specific training samples (i.e., documents) or document fragments in the corpora. The utility of sample-based explanations for LLMs is multifaceted. First, tracing back the predictions of LLM to the training samples can provide evidence for the generation results, which facilitates model debugging in cases of errors and increases the trustworthiness of the model from users when the outcomes are accurate. Second, it can also help researchers understand how LLMs generalize from training samples. If the outputs of LLMs can be traced back to exact subsequences directly spliced from the training data, it might suggest that the LLM is simply memorizing the data.
4.1. Literature Review of Sample-based Explanation
4.1.1. Influence Function-based Methods
@@
One strategy to quantify the influence of a training sample $z_{i}$ in the dataset $\mathcal{D}_{train}$ to a test sample $z$ is through the influence function (Koh and Liang, 2017; Han et al., 2020), which measures the change of the prediction loss $\mathcal{L}\left(z,\theta\right)$ for $z$, when the training sample $z_{i}$ is removed from the dataset:
(8) $\mathcal{I}(z_{i},z)=-\nabla_{\theta}\mathcal{L}(z,\hat{\theta})^{\top}\mathbf{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\mathcal{L}(z_{i},\hat{\theta}),$
@@
where $\nabla_{\theta}\mathcal{L}(z,\hat{\theta})$ is the gradient of loss $\mathcal{L}$ on $z$, and $\mathbf{H}_{\hat{\theta}}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}^{2}\mathcal{L}(z_{i},\hat{\theta})$ is the Hessian matrix. To improve efficiency, Koh and Liang (2017) adopt an iterative approximation process to calculate the Hessian-Vector Product (HVP) in Eq. (8), where the memory complexity can be reduced to $\mathcal{O}(P)$ and time complexity to $\mathcal{O}(NPr)$ ($r$ is the number of iterations). To further reduce the complexity, (Pruthi et al., 2020) propose TracIn, which measures the influence by calculating the total reduction of loss whenever $z_{i}$ is included in the minibatch during model training as follows:
@@
(9) $\mathcal{I}_{\operatorname{TracIn}}\left(z_{i},z\right)=\sum_{t:z_{i}\in\mathcal{B}_{t}}\mathcal{L}\left(z,\theta_{t}\right)-\mathcal{L}\left(z,\theta_{t+1}\right)\approx\frac{1}{|B_{t}|}\sum_{t:z_{i}\in B_{t}}\eta_{t}\nabla_{\theta}\mathcal{L}\left(z_{i},\theta_{t}\right)\cdot\nabla_{\theta}\mathcal{L}\left(z,\theta_{t}\right),$
@@
where $\mathcal{B}_{t}$ is the $t$-th mini-batch, $\theta_{t}$ is the parameter at the $t$-th step, $\eta_{t}$ is the step size. TracIn only leverages gradients, which substantially improves the efficiency. In addition, Schioppa et al. (2022) propose to use Alnordi iteration (Arnoldi, 1951) to find the dominant eigenvalues and eigenvectors of $\mathbf{H}_{\hat{\theta}}$ on randomly sampled subsets $\mathcal{D}_{sub}$, with $|\mathcal{D}_{sub}|\ll|\mathcal{D}_{train}|$, where the diagonalized Hessian can be cheaply cached and inverted. Observing that finding the most influential training sample on $z$ needs to iterate overall $N$ training samples, (Guo et al., 2021) propose to use fast KNN to pre-filter a small subset of influence-worthy data points from $\mathcal{D}_{train}$ as candidates to explain small pretrained language models, whereas (Han and Tsvetkov, 2022) propose to find a small subset $\mathcal{D}_{sub}\subset\mathcal{D}_{train}$ whose gradient is the most similar to the downstream task examples. (Grosse et al., 2023) propose to use the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) approximation to scale influence functions to 52B LLMs. For adaptation, only influences mediated by the feed-forward networks (FFNs) are considered. Based on the assumption that weights from different FFN layers are independent, the EK-FAC approximated influence is finally formulated as the sum of influences mediated by each layer as follows:
(10) $\mathcal{I}_{\text{EKFAC}}(z_{i},z)=\sum_{l}\nabla_{\theta^{(l)}}\mathcal{L}(z,\hat{\theta})^{\top}(\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}+\lambda^{(l)}\mathbf{I})^{-1}\nabla_{\theta^{(l)}}\mathcal{L}(z_{i},\hat{\theta}),$
@@
where $\theta^{(l)}$ denotes the weights of the $l$-th MLP layer, and $\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}$ is the EK-FAC approximated Gauss-Newton Hessian for $\theta^{(l)}$. Since the inversion of $L$ small ${K_{l}\times K_{l}}$ matrices (i.e., $\mathcal{O}(L\times K_{l}^{3})$) is substantially more efficient than the inversion of a large $L{K_{l}\times LK_{l}}$ matrix (i.e., $\mathcal{O}{((LK_{l}})^{3})$), $\mathcal{I}_{\text{EKFAC}}$ can be adaptable to very large models.
Recently, influence functions have also been used to explain and improve the in-context learning ability of LLMs. Based on the finding  (Von Oswald et al., 2023) that during in-context learning, LLMs can be viewed as implicitly ”learning” an internal kernelized least square surrogate objective on the few-shot examples included in the prompt, Zhou et al. (2024) propose to use this hypothesized objective as the loss function to calculate the influence of each in-context sample on the LLM generation, and based on which select and re-rank the samples. They show that the in-context learning ability of LLMs can be improved after the sample selection and demonstration order re-ranking.
4.1.2. Embedding-based Methods
@@
Another strategy for sample-based explanation leverages hidden representations within the transformer architecture, which is recognized for encoding high-level semantics from textual data, to calculate the semantic similarity between $z$ and $z_{i}$. The similarity can also be used to measure the influence of $z_{i}$ on $z$ as explanations (Rajani et al., 2019). Specifically, (Akyurek et al., 2022) propose to represent the training sample $z_{i}$ and test sample $z$ by concatenating the input and output as $z^{cat}_{i}=[x_{i}||y_{i}]$, $z^{cat}=[x||y]$. The concatenation is feasible for generation tasks where the output $y$ lies in the same token sequence space as the input prompt $x$. The similarity between $z_{i}$ and $z$ can then be calculated as:
(11) $\mathcal{I}_{\text{emb}}\left(z_{i},z\right)=\frac{f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\cdot f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)}{\left\|f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\right\|\left\|f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)\right\|},$
@@
where $f^{(l)}_{\hat{\theta}}$ is the sub-network that outputs the $l$-th layer intermediate activation of $f_{\hat{\theta}}$. The Eq. (11) has a similar form as the vanilla influence function defined in Eq. (8) as well as its TracIn alternative defined in Eq. (9), which assigns a score $\mathcal{I}$ for the explainee $z$ for each training sample $z_{i}$ in the dataset $\mathcal{D}_{train}$ as the explanation confidence of the sample $z_{i}$.
4.2. Case Study: EK-FAC-based Influence Estimation
@@
This case study implements the EK-FAC-based influence function (Grosse et al., 2023) and verifies its scalability and effectiveness on LLMs with billions of parameters, including GPT2-1.5B (Radford et al., 2019), LLaMA2-7B (Touvron et al., 2023b), Mistral-7B (Jiang et al., 2023), and LLaMA2-13B.
4.2.1. Experimental Design
@@
We use the SciFact dataset (Wadden et al., 2020) as the corpora, which contains the abstract of 5,183 papers from basic science and medicine. The LLMs are obtained by finetuning the pretrained LLMs for 20,000 iterations, where AdamW is used as the optimizer (Loshchilov and Hutter, 2018). Then, we use 500 samples from the corpora to estimate the (i) uncentered covariance matrices of the activations and pre-activation pseudo-gradients $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$, and (ii) the variances of the projected pseudo-gradient $\boldsymbol{\Lambda}^{(l)}$ for each selected dense layer $l$, and cache them on the hard disk (details see Eqs. (16) and (20) in (Grosse et al., 2023)). We select the c_fc layer for GPT2-1.5B, and gate_proj layer for LLaMA2-7B, Mistral-7B, and LLaMA2-13B. For evaluation, we randomly select 200 samples to construct the test set, which we name SciFact-Inf. Specifically, for the $j$-th sample $z_{j}=x_{j}$, we use the first three sentences in $x_{j}$, i.e., $\hat{x}_{j}$, to generate a completion $\hat{y}_{j}$ with the finetuned LLM. Ideally, the $j$-th training sample $z_{j}$ itself should be the most influential sample w.r.t. the generation of $\hat{y}_{j}$.
4.2.2. Results and Analysis